Solve the substitution method.
4x+6y=0
-9x+y=29
solution ordered pair:( , )
or
There is no solution?

Answer by Maths68(1474)

4x+6y=0........(1)
-9x+y=29........(2)
Take (1)
4x+6y=0
4x=-6y
4x/4=-6y/4
x=-3y/2
Substitute the value of x in (2)
-9x+y=29........(2)
-9(-3y/2)+y=29
27y/2+y=29
Multiply by 2 both sides of above equation
(2)(27y/2+y)=(29)(2)
27y+2y=58
29y=58
29y/29=58/29
y=2
Put the value of y in (1)
4x+6y=0........(1)
4x+6(2)=0
4x+12=0
4x=-12
4x/4=-12/4
x=-3
x,y = -3, 2

Check
======
Put the values of x and y in (1)
4x+6y=0........(1)
4(-3)+6(2)=0
-12+12=0
0=0
Put the values of x and y in (1)
-9x+y=29........(2)
-9(-3)+2=29
27+2=29
29=29